# DeepAgent: A General Reasoning Agent with Scalable Toolsets

### v2

@@ Abstract.
Large reasoning models have demonstrated strong problem-solving abilities, yet real-world tasks often require external tools and long-horizon interactions. Existing agent frameworks typically follow predefined workflows, which limit autonomous and global task completion. In this paper, we introduce DeepAgent, an end-to-end deep reasoning agent that performs autonomous thinking, tool discovery, and action execution within a single, coherent reasoning process. To address the challenges of long-horizon interactions, particularly the context length explosion from multiple tool calls and the accumulation of interaction history, we introduce an autonomous memory folding mechanism that compresses past interactions into structured episodic, working, and tool memories, reducing error accumulation while preserving critical information. To teach general-purpose tool use efficiently and stably, we develop an end-to-end reinforcement learning strategy, namely ToolPO, that leverages LLM-simulated APIs and applies tool-call advantage attribution to assign fine-grained credit to the tool invocation tokens. Extensive experiments on eight benchmarks, including general tool-use tasks (ToolBench, API-Bank, TMDB, Spotify, ToolHop) and downstream applications (ALFWorld, WebShop, GAIA, HLE), demonstrate that DeepAgent consistently outperforms baselines across both labeled-tool and open-set tool retrieval scenarios. This work takes a step toward more general and capable agents for real‑world applications. The code and demo are available at https://github.com/RUC-NLPIR/DeepAgent.
@@
* Work done during internship at Xiaohongshu.
$\dagger$ Corresponding author.
@@ 1. Introduction
Recently, the advent of Large Reasoning Models (LRMs) has demonstrated the capability to solve complex problems in domains like mathematics, programming, and scientific reasoning through a step-by-step “slow thinking” process (Chen et al., 2025a; Li et al., 2025e; Wang et al., 2025c). However, many real-world tasks necessitate the use of external tools for their completion. Recent approaches (e.g., Search-o1 (Li et al., 2025a), DeepResearcher (Zheng et al., 2025), and ToRL (Li et al., 2025d)) integrate tool use into reasoning, but typically rely on a small, fixed tool set such as search, browsing, and coding (Figure 2(b)), limiting their generality.
@@
To empower DeepAgent to thoroughly and robustly explore new tools and navigate complex environments during long-horizon interactions, we equip it with memory management capabilities. We introduce an Autonomous Memory Folding strategy that allows DeepAgent to consolidate its previous thoughts and interaction history into a structured memory schema at any point during its thinking before resuming the agentic reasoning process. This mechanism not only saves tokens and enhances reasoning efficiency over extended interactions but also provides the agent an opportunity to “take a breath”, preventing it from becoming trapped in wrong exploration paths and enabling it to reconsider its strategy. To mitigate information loss during folding, we introduce a brain-inspired memory architecture comprising episodic memory, working memory, and tool memory, all structured with an agent-usable data schema to ensure the stability and utility of the folded memory.
@@ 4.1. Tasks and Datasets
General Tool-Use. These benchmarks encompass a broad range of distinct tools, scaling from tens to over ten thousand, making them ideal for evaluating the scalability of different approaches. We utilize four representative scenarios: ToolBench (Qin et al., 2024b), based on over 16,000 real-world APIs, for which we use the G3 subset requiring multi-step, multi-tool calls; API-Bank (Li et al., 2023), which includes 314 human-annotated dialogues with 73 APIs and 753 API calls, to assess planning, retrieval, and calling capabilities; RestBench (Song et al., 2023), comprising scenarios from the TMDB movie database (54 tools, avg. 2.3 calls/question) and the Spotify music player (40 tools, avg. 2.6 calls/question) to simulate typical REST applications; and ToolHop (Ye et al., 2025), a multi-hop reasoning dataset with 3,912 locally executable tools that necessitate 3 to 7 sequential tool calls per task. For these tasks, we adopt two settings: given ground-truth tools and given entire toolsets with tool retrieval capabilities.
@@ 5.1. Main Results on General Tool Usage Tasks
Table 1 presents the results on general tool usage, leading to several key observations. (1) DeepAgent’s End-to-End Reasoning Surpasses Workflow-Based Methods. DeepAgent’s holistic agentic process consistently outperforms rigid, predefined workflows. For instance, on labeled-tool tasks, DeepAgent-32B-RL achieves success rates of 89.0% on TMDB and 75.4% on Spotify, substantially exceeding the strongest 32B baseline scores of 55.0% and 52.6%, respectively. This underscores the benefit of a holistic agentic process over rigid, predefined action cycles. (2) DeepAgent Maintains Robustness in Open-Set Scenarios. This advantage is more pronounced in open-set scenarios where dynamic tool discovery is critical. On ToolBench and ToolHop, DeepAgent-32B-RL achieves success rates of 64.0% and 40.6%, respectively, far exceeding the top baseline scores of 54.0% and 29.0%. This demonstrates that DeepAgent’s strategy of dynamically discovering tools as needed within the reasoning process is far more robust and scalable in realistic open-set scenarios. (3) ToolPO Training Further Improves Tool-Usage Capabilities. The proposed ToolPO RL strategy provides significant further gains. The trained DeepAgent-32B-RL model consistently improves upon its base version, boosting success rates on ToolBench by up to 6.0% and on Spotify (labeled) by 5.2%. This validates the effectiveness of the ToolPO strategy, which uses an LLM-based tool simulator and fine-grained advantage attribution.
@@ 5.2. Main Results on Downstream Applications
Table 2 shows the results on downstream applications, which require agents to handle long-horizon interactions in complex environments. (1) The autonomous reasoning paradigm generally outperforms workflow-based methods. On complex application tasks, methods that integrate tool usage into continuous reasoning consistently outperform rigid, predefined workflows. On GAIA, both DeepAgent-32B-Base (46.7) and HiRA (42.5) significantly exceed the best workflow-based method CodeAct (34.5). Similarly, on WebShop, DeepAgent-32B-Base (32.0) substantially surpasses CodeAct (18.0). This demonstrates that long-horizon interaction tasks require deep agentic reasoning capabilities to achieve superior task accomplishments. (2) DeepAgent demonstrates superior performance across various application tasks. DeepAgent achieves state-of-the-art performance among 32B models. On GAIA, DeepAgent-32B-RL scores 53.3 vs. HiRA’s 42.5, and on ALFWorld reaches 91.8% vs. HiRA’s 84.3%. This stems from DeepAgent’s seamless integration of actions into coherent reasoning, enabling end-to-end execution with autonomous memory folding, which is advantages unavailable to workflow-constrained methods. (3) ToolPO training further improves performance on downstream applications. ToolPO training yields consistent gains over the base model. DeepAgent-32B-RL improves GAIA scores from 46.7 to 53.3 (+6.6) and ALFWorld success rates from 88.1% to 91.8% (+3.7), demonstrating that ToolPO effectively enhances reasoning and tool usage capabilities for complex task completion.
@@ 6. Conclusion
In this work, we introduce DeepAgent, an end-to-end reasoning agent that unifies thinking, tool discovery, and execution into a single, coherent agentic reasoning process. To enable robust long-horizon interaction, we propose an autonomous memory folding mechanism that compresses interaction history into a structured memory, allowing the agent to ”take a breath” and reconsider its strategy. We also introduce ToolPO, an end-to-end RL method that leverages LLM simulated APIs for stable training and fine-grained advantage attribution for precise credit assignment to tool invocations. Extensive experiments on general tool-use and downstream applications demonstrate that DeepAgent significantly outperforms various baseline agents, particularly in open-set scenarios requiring dynamic tool discovery over scalable toolsets. This work opens new avenues for creating more general and capable agents, with future research directions including developing advanced memory management, multimodal (or even omni-modal) capabilities, and human-agent collaborative reasoning.

### v3

@@ Abstract.
Large reasoning models have demonstrated strong problem-solving abilities, yet real-world tasks often require external tools and long-horizon interactions. Existing agent frameworks typically follow predefined workflows, which limit autonomous and global task completion. In this paper, we introduce DeepAgent, an end-to-end deep reasoning agent that performs autonomous thinking, tool discovery, and action execution within a single, coherent reasoning process. To manage long-horizon interactions, we introduce an autonomous memory folding mechanism that compresses past interactions into structured episodic, working, and tool memories, reducing error accumulation while preserving critical information. To teach general-purpose tool use efficiently and stably, we develop an end-to-end reinforcement learning strategy, namely ToolPO, that leverages LLM-simulated APIs and applies tool-call advantage attribution to assign fine-grained credit to the tool invocation tokens. Extensive experiments on eight benchmarks, including general tool-use tasks (ToolBench, API-Bank, TMDB, Spotify, ToolHop) and downstream applications (ALFWorld, WebShop, GAIA, HLE), demonstrate that DeepAgent consistently outperforms baselines across both labeled-tool and open-set tool retrieval scenarios. The code and demo are available at https://github.com/RUC-NLPIR/DeepAgent.
@@ 1. Introduction
Recently, the advent of Large Reasoning Models (LRMs) has demonstrated the capability to solve complex problems in domains like mathematics, programming, and scientific reasoning through a step-by-step “slow thinking” process (Chen et al., 2025a; Li et al., 2025e; Wang et al., 2025c). However, many real-world tasks necessitate the use of external tools for their completion. Recent approaches integrate tool use into reasoning (Li et al., 2025a; Zheng et al., 2025; Li et al., 2025d), but typically rely on a small, fixed tool set such as search, browsing, and coding (Figure 2(b)), limiting their generality.
@@
To facilitate robust exploration in long-horizon environments, we equip DeepAgent with Autonomous Memory Folding. This strategy allows the agent to dynamically consolidate its reasoning process and interaction history into a structured memory schema. Beyond reducing token overhead, this mechanism enables the agent to “take a breath”—pausing to reconsider strategies and avoid erroneous paths. To minimize information loss during consolidation, we introduce a brain-inspired memory architecture comprising episodic, working, and tool memory, all structured with an agent-usable data schema to ensure the stability and utility of the folded memory.
@@ 4.1. Tasks and Datasets
General Tool-Use. These benchmarks cover toolsets from tens to >10>10k tools, and thus stress scalability. They evaluate core capabilities for general tool use, including tool planning, tool retrieval, and accurate multi-step tool calling. We use ToolBench (Qin et al., 2024b) (16k+ real-world APIs; G3 subset with multi-step/multi-tool calls), API-Bank (Li et al., 2023) (314 dialogues; 73 APIs; 753 calls), RestBench (Song et al., 2023) (TMDB: 54 tools, 2.3 calls/question; Spotify: 40 tools, 2.6 calls/question), and ToolHop (Ye et al., 2025) (3,912 executable tools; 3–7 calls/task). We evaluate two settings: provided ground-truth tools and open-set tool retrieval from the full toolset.
@@ 5.1. Main Results on General Tool Usage Tasks
Table 1 summarizes results on general tool-use tasks and yields three observations. (1) DeepAgent’s End-to-End Reasoning Surpasses Workflow-Based Methods. DeepAgent consistently outperforms workflow-based agents. On labeled-tool tasks, DeepAgent-32B-RL reaches 89.0% on TMDB and 75.4% on Spotify, exceeding the best 32B baselines (55.0% and 52.6%). This highlights the advantage of end-to-end agentic reasoning over rigid, predefined action loops. (2) DeepAgent Maintains Robustness in Open-Set Scenarios. Gains are larger in open-set settings where tool discovery is required: on ToolBench and ToolHop, DeepAgent-32B-RL achieves 64.0% and 40.6%, surpassing the best baselines (54.0% and 29.0%). This suggests that on-demand tool discovery within the reasoning process is both more robust and more scalable in realistic open-set tool environments. (3) ToolPO Training Further Improves Tool-Usage Capabilities. ToolPO yields consistent improvements over the base model, increasing ToolBench success by up to 6.0% and Spotify (labeled) by 5.2%. These gains indicate that our RL training better aligns intermediate tool calls with end-task success.
@@ 5.2. Main Results on Downstream Applications
Table 2 reports the downstream results that require long-horizon interaction and more complex environment dynamics. (1) The autonomous reasoning paradigm generally outperforms the workflow-based methods. Methods that integrate tool use into continuous reasoning outperform workflow-based agents. On GAIA, DeepAgent-32B-Base (46.7) and HiRA (42.5) exceed the best workflow baseline CodeAct (34.5); on WebShop, DeepAgent-32B-Base scores 32.0 vs. 18.0. This supports that long-horizon tasks benefit from flexible, integrated reasoning-and-action rather than fixed workflows. (2) DeepAgent demonstrates superior performance across various application tasks. DeepAgent achieves the best performance among 32B models: 53.3 on GAIA (vs. 42.5 for HiRA) and 91.8% on ALFWorld (vs. 84.3). We attribute this to DeepAgent’s coherent reasoning process and its support for robust long-horizon interaction (e.g., autonomous memory folding). (3) ToolPO training further improves performance on downstream applications. ToolPO further improves downstream performance: GAIA 46.7 $\rightarrow$ 53.3 (+6.6) and ALFWorld 88.1% $\rightarrow$ 91.8% (+3.7). This shows the tool-use improvements learned by ToolPO transfer to interactive downstream settings.
@@ 6. Conclusion
In this work, we introduce DeepAgent, an end-to-end reasoning agent that unifies thinking, tool discovery, and execution into a single, coherent agentic reasoning process. To enable robust long-horizon interaction, we propose an autonomous memory folding mechanism that compresses interaction history into a structured memory, allowing the agent to ”take a breath” and reconsider its strategy. We also introduce ToolPO, an end-to-end RL method that leverages LLM simulated APIs for stable training and fine-grained advantage attribution for precise credit assignment to tool invocations. Extensive experiments on general tool-use and downstream applications demonstrate that DeepAgent significantly outperforms various baseline agents, particularly in open-set scenarios requiring dynamic tool discovery over scalable toolsets. This work opens new avenues for developing more general and scalable LLM agents for broader real-world applications.